A sinusoidal transverse wave is travelling along a
Answer #1

a). From the plot given we can see that maximum amplitude is,

  A= 5(1cm)= 5cm (as 1 unit on y axis is equal to 1cm)

b). Wavelength is defined as the distance between successive crests or successive troughs.

Then from the plot, we get wavelength as, \lambda= 4(10cm)= 40cm

(as 1 unit on x axis is equal to 10cm)

c). We know that wave speed of strings, v= \sqrt{\frac{T}{\mu}}

where T= Tension in the string= 3.3 N

\mu= linear density= 44g/m= 0.044 kg/m

then wave speed will be v= \sqrt{\frac{3.3}{0.044}}= \sqrt{68.181}= 8.257 m/s

d). We know that if "f" is the frequency of the wave then we'll have

v= f\lambda

and Time Period, T= \frac{1}{f}

then we'll have T= \frac{\lambda}{v}

using all calculated values in above, T= \frac{0.40}{8.257}= 0.0484 s

e). Maximum Transverse Speed is given by, v_{max}= \omega \times A= ( \frac{2 \pi }{T})A

  v_{max}= ( \frac{2 \times 3.14 }{0.0484})(0.05)= 6.487 m/s

f). If the given equation depicts the given wave and plot, then wavenumber(k) is given by

  k= \frac{2 \pi }{\lambda}= \frac{2 \times 3.14 }{0.4}= 15.7 m^{-1}

g). And the angular speed is given by \omega= \frac{2 \pi }{T}= \frac{2 \times 3.14 }{0.0484}= 129.75s^{-1}

h). Using above values we can write the equation as

Y(x,t)= y_msin(kx-\omega t+\phi )

Now since wave travels in negative x direction, hence

  Y(x,t)= Asin(-15.7x-129.75t+\phi )

  Y(x,t)= Asin(15.7x+129.75t+\phi )

Then using the value of amplitude and value of displacement at time t=0s,

  0.04= 0.05sin(15.7(0)+129.75(0)+\phi )

  0.8= sin(\phi )

  \phi= sin^{-1}(0.8)= 53.130^{\circ}

i). The sign will become positive in front of \omega as shown above.